# Automated Composition of Agents: A Knapsack Approach for Agentic Component Selection

### v1

@@ 1 Introduction
The design of effective agentic systems sits at the frontier of artificial intelligence research, promising autonomous agents capable of sophisticated reasoning, tool manipulation, and collaborative problem-solving. Yet as the ecosystem of AI components expands, with proliferating models, APIs, and specialized agents, a critical bottleneck emerges: the paradox of choice [23]. While modular reuse offers clear advantages over building systems from scratch, developers face a combinatorial explosion of possible configurations, each with hidden constraints and unpredictable interactions. Traditional approaches relying on manual curation or metadata-based retrieval [29] struggle with three fundamental limitations: opaque capability descriptions that rarely match real-world performance, myopic selection criteria that ignore cost-utility trade-offs, and static architectures that break down when requirements evolve.
@@
    A formalization of agent composition as a constrained optimization problem that jointly considers capability, cost, and compatibility, bridging gaps between modular AI design [10] and operations research [41].
@@
Design of machine learning systems is critical as poor organization leads to unstable dependencies, pipeline jungles, and many other types of hidden technical debt [28]. Our work is a step towards optimizing this process to reduce these issues and maintain a well-composed agentic system. As we explore in subsequent sections, our approach is applicable to composing both single agents and multi-agent systems.
@@ 2 Related Work
The novelty of AI agents stems from the integration of APIs with LLMs [27, 24]. Since tools are a critical component of agents, prior works have focused on retrieval approaches to select the most appropriate tools. Qin et al. [25] collect a large dataset of tools based on RapidAPIs and train a BERT-based tool retriever. Shi et al. [29] note that tool retrieval is a difficult problem as commonly used retrievers fail to capture user intent and select the most relevant tools. RAG-MCP [9] combines RAG with the Model Context Protocol (MCP) to improve how LLMs select external tools. Recently, more work has emphasized the vulnerability of agents due to poor selection of tools [6, 22].
@@
Hu et al. [10] introduce the problem of automated design of agentic systems (ADAS), where the goal is to “automatically create powerful agentic system designs, including inventing novel building blocks and/or combining them in new ways”. Conceptually, the building blocks of ADAS includes novel prompts, tool use, and workflows. In our paper, we focus on a subset of ADAS, which is automated composition of agentic systems. Our setting assumes that the underlying building blocks exist and the algorithmic challenge is to select the most optimal subset of those components. Common solutions to this problem include tool retrieval and agent selection. For agent selection, existing works treat the problem as graph optimization where nodes are the individual agents and edges represent communication channels. DyLAN [17] introduce the agent selection problem and train a feed-forward network to optimize the selection. AgentPrune [40] extend selection to target improving communication redundancy. Multi-agent Architecture Search [39] optimizes an agentic supernet and then will sample a multi-agent architecture in a query-dependent manner. Wu et al. [35] jointly optimize agent prompt and tool descriptions to improve workflow efficiency. The component selection problem also relates to Distributed Constraint Optimization Problems (DCOP) [8], where traditional agents assign values to variables to minimize a global cost function subject to constraints. DCOP techniques have been applied to service selection and composition problems, particularly for QoS-aware web service composition [4].
@@
The knapsack problem, a classic optimization challenge, has been extensively studied in algorithmic research [19, 7, 14]. Offline algorithms, such as dynamic programming [2] and branch-and-bound methods [18], assume complete knowledge of all items in advance, enabling optimal solutions for static inputs. In contrast, online knapsack algorithms process items sequentially without future information. A commonly known approach is the ZCL algorithm [41] which proposes a threshold for the value-to-weight ratio and is dynamically configured based on the capacity of the knapsack. Other theoretically competitive approaches further extend the ZCL algorithm [11, 16].
@@ 3 Problem Definition
Hu et al. [10] introduce the problem of automated design of agentic systems (ADAS), where the goal is to automatically build and configure the components of an agentic system. Rather than manually engineering agent architectures, ADAS aims to automate the entire procedure through algorithmic discovery. Components of agents include tools, models, prompts, and workflows. Our work addresses a critical subproblem within this vision: automated agent composition, which focuses not on inventing new building blocks from scratch, but on optimally selecting and combining existing components from inventories to meet specific task requirements.
@@ 4 Composer Agents: From Semantic Retrieval to Knapsack Selection
Given task $\tau$ and its description $x$, the retrieval composer will first generate a list of skills $m\in\mathcal{M}$, each with a name and a description. The retrieval composer then queries the inventory with the skill name and description to find the most relevant component. The selected subset $\mathcal{S}$ will be the top-1 retrieval results based on the skills. Note that semantic matching has been used in prior work in agent discovery [5] and more generally in service discovery [20]. Therefore, we include a retrieval-only baseline in our experiments for fair comparison.
@@
Offline Knapsack Composer: The above composers do not take budget $B$ into account when making the selection. To find a solution $\mathcal{S}$ with success rate $p_{\tau}(S)$ close to the optimal success rate $p_{\tau}(S^{*})$ and following the budget constraint, we apply solutions of the Knapsack problem. First, we generate skills based on task description $x$ (Appendix A.1) and retrieve top-$K$ components for each skill. Next, we assign the value of each component based on the similarity score (SIM) used in retrieval. Finally, we use linear programming to find the optimal solution for this multiple-choice knapsack problem [31]:
@@
In our problem, the value of the component is not known until it is tested. Therefore, we use the ZCL algorithm [41], a commonly used online knapsack algorithm in the literature. The algorithm assumes there is a lower and upper bound, $L$ and $U$, on the value-to-cost ratio. The solution is theoretically proven to be $\ln(U/L)+1$-competitive. The algorithm sets a dynamic threshold based on the capacity filled in the knapsack and only accepts an incoming item if its value-to-cost ratio exceeds the threshold $\Psi$. To compute the value, the composer agent has to not only generate the skills but also a set of test questions for each skill (Algorithm 1). The composer agent then uses those questions to evaluate the component. Table 1 shows the test queries that are generated and then used to test tools during the sandboxing trials. Appendix A.2 goes over the prompt and details on how the component is judged. The component’s value is then based on this assessment.
If a component is successfully equipped, then the composer agent will move forward to the next skill and iterate over the top-K components until it finds a successful one. Note to further optimize upon the runtime of Algorithm 1, we make modifications such as: 1) stopping sandboxing early if component breaks and preventing it from being tested again, 2) once a skill is covered, we do not assess that skill again for other candidate tools. We cover these shortcut optimizations in Appendix A.3.
@@
Initialize $\mathcal{S}\leftarrow\emptyset$, remaining budget $\hat{B}\leftarrow B$;
Estimate value-to-cost bounds $U\leftarrow\max_{a\in\mathcal{A}}\frac{v_{a}}{c_{a}}$, $L\leftarrow\min_{a\in\mathcal{A}}\frac{v_{a}}{c_{a}}$;
@@
// Returns $\{(m_{j},d_{j},Q_{j})\}$ (name, desc, test questions)
@@
 $\mathcal{T}_{j}\leftarrow\textsc{TopKComponents}(\mathcal{A},d_{j},K)$ ;
 // Top $K$ by $\textsc{SIM}(d_{a},d_{j})$
@@
    $score\leftarrow 0$;
    foreach question $q\in Q_{j}$ do
       $response\leftarrow\textsc{Sandbox}(a_{i},q)$ ;
       // Sandbox component
       $judgement\leftarrow\textsc{Judge}(response,q,m_{j})$ ;
       // Returns 0/1
       $score\leftarrow score+judgement$
      end foreach
    $v_{i}\leftarrow\frac{score}{|M_{j}|}\cdot U$ ;
    // Normalized value estimation
    $\rho_{i}\leftarrow\frac{v_{i}}{c_{i}}$ ;
@@
    $\Psi\leftarrow(\frac{Ue}{L})^{\hat{B}}(\frac{L}{e})$ ;
    // ZCL threshold
    if $\rho_{i}\geq\Psi$ and $c_{i}\leq\hat{B}$ then
       $\mathcal{S}\leftarrow\mathcal{S}\cup\{a_{i}\}$;
       $\hat{B}\leftarrow\hat{B}-c_{i}$;
         break ;
       // Go to next skill
@@ 5.1 Composing Single Agents
For single-agent experiments, the inventory $\mathcal{A}$ consists of 120 tools. For this tool inventory, we first collect tools that are actual APIs and can be easily found on Langchain [15], like arXiv, PubMed, SemanticScholar, web search, etc. For the rest of the tools, we collect a small subset from the largest available tool retrieval benchmark, ToolRet [29]. We discover that many tools in this benchmark are not readily available to use and very specialized for one usecase. This is likely because the benchmark is used to test retrieval accuracy. For more details on the construction of the tool inventory, refer to Appendix A.4. For pricing, we estimate the cost of a tool to consist of the input tokens (from the tool schemas) during inference and cost of the API call. The pricing for about 5K calls to the agent for a free API tool is around \$3. For the paid APIs, we then estimate the tools to be about \$5 and \$8. See Appendix A.5 for more details.
@@
We fix the agent workflow to CodeAct [33], where the agent is prompted to generate code in order to invoke tools rather than in standardized JSON format [38]. We choose CodeAct as it has shown to outperform other tool-calling frameworks. We experiment with both Claude 3.5 Sonnet (claude-3-5-sonnet-20241022), Claude 3.5 Haiku (claude-3-5-haiku-20241022[3], and Claude 3.7 Sonnet (claude-3-7-sonnet-20250219). In one experiment, we will use the same model for the composer agent and the candidate agent that is being evaluated. For embedding model, we use BGE-Large-English embeddings (bge-large-en-v1.5[37].
@@
For single-agent experiments, we evaluate on GAIA, SimpleQA, and MedQA. GAIA [21], which stands for General AI Assistants, is a comprehensive evaluation framework designed to assess the capabilities of AI systems in handling real-world scenarios. It focuses on testing fundamental abilities such as reasoning, multi-modal understanding, web browsing, and tool-use proficiency. Unlike traditional benchmarks that challenge AI with tasks difficult for humans, GAIA poses questions that are conceptually simple for humans but pose significant challenges for advanced AI models. SimpleQA [34] is a factuality evaluation framework developed by OpenAI to measure the ability of language models to answer short, fact-seeking questions accurately. The questions in SimpleQA are crafted to have a single, indisputable answer, making it easier to evaluate the factual correctness of model responses. MedQA [12] is a comprehensive evaluation framework designed to assess the clinical knowledge of language models in healthcare settings. Questions are derived from the United States Medical License Exams and aims to evaluate the ability of agentic systems to apply medical knowledge in practical scenarios. Note that we reuse the “smolagents” version of GAIA and SimpleQA, which is an active leaderboard for LLM agents hosted on Huggingface [26].
@@
For retrieval, we set $K$ to 10. For question generation, we set the number of test questions per skill to 2. During judgement, we ask the composer to judge whether the tool was useful for the agent to answer the question. Additionally, we include an approach where we conduct tool selection through online knapsack and then conduct an additional round of prompt optimization using AvaTaR [36]. The feedback for prompt optimization is derived from the tool sandboxing trajectories when the composer has to test the CodeAct agent with the tool. Therefore, we already have existing twelve trajectories per tool for prompt optimization, so we can directly apply AvaTaR. This helps refine the agent’s prompt so that it has a better understanding of when to invoke certain tools.
@@ Results
Figure 3 shows the results from the Claude 3.5 Sonnet experiments where we plot success rate against the budget spent on tools. Following the accuracy-cost analysis done by Kapoor et al. [13], we plot the pareto frontier. Approaches that are on the pareto frontier signify that there is no other method that outperforms them simultaneously in both dimensions. Overall, we observe the highest performing approach to be online knapsack with AvaTaR optimization (\$30 budget constraint). On all three datasets, its success rate is higher than identity but cost is much lower. Similarly, online knapsack without AvaTaR optimization shows to be on the pareto frontier for GAIA and MedQA, and has relatively higher success rate than the other approaches. Retrieval-only approaches (retrieval and offline knapsack) tend to fare worse across all three datasets, which validates previous findings that only using retrieval for tool discovery is insufficient [29].

### v2

@@ 1 Introduction
The design of effective agentic systems sits at the frontier of artificial intelligence research, promising autonomous agents capable of sophisticated reasoning, tool manipulation, and collaborative problem-solving. Yet as the ecosystem of AI components expands, with proliferating models, APIs, and specialized agents, a critical bottleneck emerges: the paradox of choice [oulasvirta2009more]. While modular reuse offers clear advantages over building systems from scratch, developers face a combinatorial explosion of possible configurations, each with hidden constraints and unpredictable interactions. Traditional approaches relying on manual curation or metadata-based retrieval [shi2025retrieval] struggle with three fundamental limitations: opaque capability descriptions that rarely match real-world performance, myopic selection criteria that ignore cost-utility trade-offs, and static architectures that break down when requirements evolve.
@@
    A formalization of agent composition as a constrained optimization problem that jointly considers capability, cost, and compatibility, bridging gaps between modular AI design [hu2024automated] and operations research [zhou2008budget].
@@
Design of machine learning systems is critical as poor organization leads to unstable dependencies, pipeline jungles, and many other types of hidden technical debt [sculley2015hidden]. Our work is a step towards optimizing this process to reduce these issues and maintain a well-composed agentic system. As we explore in subsequent sections, our approach is applicable to composing both single agents and multi-agent systems.
@@ 2 Related Work
The novelty of AI agents stems from the integration of APIs with LLMs [schick2023toolformer, patil2024gorilla]. Since tools are a critical component of agents, prior works have focused on retrieval approaches to select the most appropriate tools. qin2024toolllm collect a large dataset of tools based on RapidAPIs and train a BERT-based tool retriever. shi2025retrieval note that tool retrieval is a difficult problem as commonly used retrievers fail to capture user intent and select the most relevant tools. RAG-MCP [rag-mcp] combines RAG with the Model Context Protocol (MCP) to improve how LLMs select external tools. Recently, more work has emphasized the vulnerability of agents due to poor selection of tools [faghih2025gaming, milev2025toolfuzz].
@@
hu2024automated introduce the problem of automated design of agentic systems (ADAS), where the goal is to “automatically create powerful agentic system designs, including inventing novel building blocks and/or combining them in new ways”. Conceptually, the building blocks of ADAS includes novel prompts, tool use, and workflows. In our paper, we focus on a subset of ADAS, which is automated composition of agentic systems. Our setting assumes that the underlying building blocks exist and the algorithmic challenge is to select the most optimal subset of those components. Common solutions to this problem include tool retrieval and agent selection. For agent selection, existing works treat the problem as graph optimization where nodes are the individual agents and edges represent communication channels. DyLAN [liu2024dynamic] introduce the agent selection problem and train a feed-forward network to optimize the selection. AgentPrune [zhang2025cut] extend selection to target improving communication redundancy. Multi-agent Architecture Search [zhang2025multiagentsearch] optimizes an agentic supernet and then will sample a multi-agent architecture in a query-dependent manner. wu2025joint jointly optimize agent prompt and tool descriptions to improve workflow efficiency. The component selection problem also relates to Distributed Constraint Optimization Problems (DCOP) [fioretto2018distributed], where traditional agents assign values to variables to minimize a global cost function subject to constraints. DCOP techniques have been applied to service selection and composition problems, particularly for QoS-aware web service composition [charif2006dynamic].
@@
The knapsack problem, a classic optimization challenge, has been extensively studied in algorithmic research [mathews1896partition, feuerman1973mathematical, karp2009reducibility]. Offline algorithms, such as dynamic programming [andonov2000unbounded] and branch-and-bound methods [martello1990knapsack], assume complete knowledge of all items in advance, enabling optimal solutions for static inputs. In contrast, online knapsack algorithms process items sequentially without future information. A commonly known approach is the ZCL algorithm [zhou2008budget] which proposes a threshold for the value-to-weight ratio and is dynamically configured based on the capacity of the knapsack. Other theoretically competitive approaches further extend the ZCL algorithm [im2021online, lechowicztime].
@@ 3 Problem Definition
hu2024automated introduce the problem of automated design of agentic systems (ADAS), where the goal is to automatically build and configure the components of an agentic system. Rather than manually engineering agent architectures, ADAS aims to automate the entire procedure through algorithmic discovery. Components of agents include tools, models, prompts, and workflows. Our work addresses a critical subproblem within this vision: automated agent composition, which focuses not on inventing new building blocks from scratch, but on optimally selecting and combining existing components from inventories to meet specific task requirements.
@@ 4 Composer Agents: From Semantic Retrieval to Knapsack Selection
Given task $\tau$ and its description $x$, the retrieval composer will first generate a list of skills $m\in\mathcal{M}$, each with a name and a description. The retrieval composer then queries the inventory with the skill name and description to find the most relevant component. The selected subset $\mathcal{S}$ will be the top-1 retrieval results based on the skills. Note that semantic matching has been used in prior work in agent discovery [cheninternet] and more generally in service discovery [meshkova2008survey]. Therefore, we include a retrieval-only baseline in our experiments for fair comparison.
@@
Offline Knapsack Composer: The above composers do not take budget $B$ into account when making the selection. To find a solution $\mathcal{S}$ with success rate $p_{\tau}(S)$ close to the optimal success rate $p_{\tau}(S^{*})$ and following the budget constraint, we apply solutions of the Knapsack problem. First, we generate skills based on task description $x$ (Appendix A.1) and retrieve top-$K$ components for each skill. Next, we assign the value of each component based on the similarity score (SIM) used in retrieval. Finally, we use linear programming to find the optimal solution for this multiple-choice knapsack problem [sinha1979multiple]:
@@
In our problem, the value of the component is not known until it is tested. Therefore, we use the ZCL algorithm [zhou2008budget], a commonly used online knapsack algorithm in the literature. The algorithm assumes there is a lower and upper bound, $L$ and $U$, on the value-to-cost ratio. The solution is theoretically proven to be $\ln(U/L)+1$-competitive. The algorithm sets a dynamic threshold based on the capacity filled in the knapsack and only accepts an incoming item if its value-to-cost ratio exceeds the threshold $\Psi$. To compute the value, the composer agent has to not only generate the skills but also a set of test questions for each skill (Algorithm 1). The composer agent then uses those questions to evaluate the component. Table 1 shows the test queries that are generated and then used to test tools during the sandboxing trials. Appendix A.2 goes over the prompt and details on how the component is judged. The component’s value is then based on this assessment.
If a component is successfully equipped, then the composer agent will move forward to the next skill and iterate over the top-K components until it finds a successful one. Note to further optimize upon the runtime of Algorithm 1, we make modifications such as: 1) stopping sandboxing early if component breaks and preventing it from being tested again, 2) once a skill is covered, we do not assess that skill again for other candidate tools. We cover these shortcut optimizations in Appendix A.3 and algorithmic details in Algorithm  2.
@@
$\mathcal{S}\leftarrow\emptyset$, $\hat{B}\leftarrow B$, $broken\leftarrow\emptyset$ ;
// $\hat{B}$: remaining budget
@@
// $\{(m_{j},d_{j},Q_{j},w_{j})\}$: name, desc, queries, importance
foreach skill $m_{j}\in\mathcal{M}$ do $\mathcal{T}_{j}\leftarrow\textsc{TopK}(\mathcal{A},d_{j},K)$;
$L\leftarrow 1/\max_{a}c_{a}$, $U\leftarrow\sum_{j}w_{j}/\min_{a}c_{a}$ ;
// Value-to-cost bounds
for $round\leftarrow 1$ to $R$ do
 $covered\leftarrow\emptyset$ ;
 // Reset each round
@@
       if $a_{i}\in\mathcal{S}\cup broken$ or $m_{j}\in covered$ or $c_{i}>\hat{B}$ then continue;
       $scores\leftarrow\textsc{Evaluate}(a_{i},\mathcal{M})$ ;
       // Returns $\{-1,0,1\}$ per skill
       // 11: helpful, 0: not helpful, 1-1: broken
       if $scores[m_{j}]<0$ then
          $broken\leftarrow broken\cup\{a_{i}\}$ ;
          // Never test again
          continue;
         end if
       $v_{i}\leftarrow\sum_{m_{k}\notin covered}w_{k}\cdot\mathbbm{1}[scores[m_{k}]=1]$ ;
       // Value from new skills
       $\rho_{i}\leftarrow v_{i}/c_{i}$ ;
@@
       $z\leftarrow(B-\hat{B})/B$, $\Psi\leftarrow(\frac{Ue}{L})^{z}\frac{L}{e}$ ;
       // Fraction spent; ZCL threshold
       if $\rho_{i}\geq\Psi$ then
          $\mathcal{S}\leftarrow\mathcal{S}\cup\{a_{i}\}$, $\hat{B}\leftarrow\hat{B}-c_{i}$, $covered\leftarrow covered\cup\{m_{k}:scores[m_{k}]=1\}$;
@@ 5.1 Composing Single Agents
For single-agent experiments, the inventory $\mathcal{A}$ consists of 120 tools. For this tool inventory, we first collect tools that are actual APIs and can be easily found on Langchain [langchain], like arXiv, PubMed, SemanticScholar, web search, etc. For the rest of the tools, we collect a small subset from the largest available tool retrieval benchmark, ToolRet [shi2025retrieval]. We discover that many tools in this benchmark are not readily available to use and very specialized for one usecase. This is likely because the benchmark is used to test retrieval accuracy. For more details on the construction of the tool inventory, refer to Appendix A.4. For pricing, we estimate the cost of a tool to consist of the input tokens (from the tool schemas) during inference and cost of the API call. The pricing for about 5K calls to the agent for a free API tool is around \$3. For the paid APIs, we then estimate the tools to be about \$5 and \$8. See Appendix A.5 for more details.
@@
We fix the agent workflow to CodeAct [wang2024executable], where the agent is prompted to generate code in order to invoke tools rather than in standardized JSON format [yaoreact]. We choose CodeAct as it has shown to outperform other tool-calling frameworks. We experiment with both Claude 3.5 Sonnet (claude-3-5-sonnet-20241022), Claude 3.5 Haiku (claude-3-5-haiku-20241022[claude35], and Claude 3.7 Sonnet (claude-3-7-sonnet-20250219). In one experiment, we will use the same model for the composer agent and the candidate agent that is being evaluated. For embedding model, we use BGE-Large-English embeddings (bge-large-en-v1.5[bge_embedding].
@@
For single-agent experiments, we evaluate on GAIA, SimpleQA, and MedQA. GAIA [mialon2023gaia], which stands for General AI Assistants, is a comprehensive evaluation framework designed to assess the capabilities of AI systems in handling real-world scenarios. It focuses on testing fundamental abilities such as reasoning, multi-modal understanding, web browsing, and tool-use proficiency. Unlike traditional benchmarks that challenge AI with tasks difficult for humans, GAIA poses questions that are conceptually simple for humans but pose significant challenges for advanced AI models. SimpleQA [wei2024measuring] is a factuality evaluation framework developed by OpenAI to measure the ability of language models to answer short, fact-seeking questions accurately. The questions in SimpleQA are crafted to have a single, indisputable answer, making it easier to evaluate the factual correctness of model responses. MedQA [jin2021disease] is a comprehensive evaluation framework designed to assess the clinical knowledge of language models in healthcare settings. Questions are derived from the United States Medical License Exams and aims to evaluate the ability of agentic systems to apply medical knowledge in practical scenarios. Note that we reuse the “smolagents” version of GAIA and SimpleQA, which is an active leaderboard for LLM agents hosted on Huggingface [smolagents].
@@
For retrieval, we set $K$ to 10. For question generation, we set the number of test questions per skill to 2. During judgement, we ask the composer to judge whether the tool was useful for the agent to answer the question. Additionally, we include an approach where we conduct tool selection through online knapsack and then conduct an additional round of prompt optimization using AvaTaR [wu2024AvaTaR]. The feedback for prompt optimization is derived from the tool sandboxing trajectories when the composer has to test the CodeAct agent with the tool. Therefore, we already have existing twelve trajectories per tool for prompt optimization, so we can directly apply AvaTaR. This helps refine the agent’s prompt so that it has a better understanding of when to invoke certain tools.
@@ Results
Figure 3 shows the results from the Claude 3.5 Sonnet experiments where we plot success rate against the budget spent on tools. Following the accuracy-cost analysis done by kapoor2024ai, we plot the pareto frontier. Approaches that are on the pareto frontier signify that there is no other method that outperforms them simultaneously in both dimensions. Overall, we observe the highest performing approach to be online knapsack with AvaTaR optimization (\$30 budget constraint). On all three datasets, its success rate is higher than identity but cost is much lower. Similarly, online knapsack without AvaTaR optimization shows to be on the pareto frontier for GAIA and MedQA, and has relatively higher success rate than the other approaches. Retrieval-only approaches (retrieval and offline knapsack) tend to fare worse across all three datasets, which validates previous findings that only using retrieval for tool discovery is insufficient [shi2025retrieval].